OpenStudy (anonymous):

w2(x + 8) + z(x + 8) = ( )(x + 8) What do I even freaking do?!

OpenStudy (anonymous):

what is the question anyway?

OpenStudy (anonymous):

Find the missing factor. I don't even know how to do this! Plato is stupid.

OpenStudy (anonymous):

I believe the missing factor is (W2+z) Whenever you have a(x+b)+c(x+b), we can simplify this as: (a+c)(x+b)

OpenStudy (mathstudent55):

The 2 is an exponent of the w, right? \(w^2(x + 8) + z(x + 8) = ( ~~~~~~~~~~)(x + 8) \)

OpenStudy (anonymous):

yes. honestly I hate plato and that's why I have to get this class done.

OpenStudy (anonymous):

Do you understand what I put? It's reversing the distributive property!

OpenStudy (mathstudent55):

If you had \(2(a) + 3(a)\), do you see that a is a common factor in 2a and 3a? Then you would be able to factor out the common factor of a leaving: \(=(2 + 3)(a)\) Using these simpler terms, do you see how 2a + 3a factors into (2 + 3)(a) ?

OpenStudy (anonymous):

HOW THE HECK WOULD YOU GO ABOUT DOING THAT?! I really can't wait to get my GED. This high school is shyt.

OpenStudy (anonymous):

honestly, I have no idea what tht even means.it just makes me want to stab myself in the juggular.

OpenStudy (anonymous):

Okay, let's take this step by step. If you have a + a, you can pull out the a, right? So that means we get a(1 + 1). This is a very simple example. Now, distribute the a back to each term a(1)+a(1) = a + a. Do you understand that so far?

OpenStudy (mathstudent55):

Your problem is similar to the 2a + 3a problem, only the terms are a little more complicated. You have: \(w^2\color{red}{(x + 8)} + z\color{red}{(x + 8)}\) Notice the common factor in red above. You can factor out the common factor, leaving: \(=\color{red}{(x + 8)}(w^2 + z) \) Now just change the order of the factors to get to your answer: \(= (w^2 + z)(x + 8)\)

OpenStudy (anonymous):

I just don't get this. \Maybe I'm stupid, but this is just... I have no idea why the Frick you would need to freaking NEED to pull out the gd a. WHY?! THE A=1? WAIT... WHAT?!

OpenStudy (mathstudent55):

You need to learn the distributive property and how to use it.

OpenStudy (anonymous):

pemdas?

OpenStudy (anonymous):

Keep in mind that we're not trying to *solve* anything. We're just trying to rearrange what we already have. So if we have 6, we can pull out factors 2 and 3 because (2)(3)=6 yeah? Similarly, if we have 4(a+1)+3(a+1) Then we can pull out the a+1 so (a+1)(4+3)

OpenStudy (mathstudent55):

Let's start with a simple example with numbers. What is 2(3 + 5) = Since we are dealing with only numbers, we can evaluate the expression. Using PEMDAS, we do the parentheses first. What is 3 + 5 = ?

OpenStudy (anonymous):

Yep ^

OpenStudy (anonymous):

8.

OpenStudy (anonymous):

I recommend watching this video by KhanAcademy. It's a very helpful video https://www.khanacademy.org/math/pre-algebra/order-of-operations/ditributive_property/v/distributive-property-example-1